A 2 kg box is sliding down a ramp with a constant velocity of 0.7 m/s. What is the net force on the box?

  

10 N

   

1.4 N

   

20 N

   

0 N
Answer #1

0 N . box is sliding with constant velocity so a=dv/dt = rate of change in velocity = 0

F=ma = 2*0 =0 N